# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v2

@@
Enhancing the precision of measurements by harnessing entanglement is a long-sought goal in the field of quantum metrology 1, 2. Yet attaining the best sensitivity allowed by quantum theory in the presence of noise is an outstanding challenge, requiring optimal probe-state generation and readout strategies 3, 4, 5, 6, 7. Neutral atom optical clocks 8, leading systems for measuring time, have shown recent progress in terms of entanglement generation 9, 10, 11, but currently lack the control capabilities to realize such schemes. Here we show universal quantum operations and ancilla-based readout for ultranarrow optical transitions of neutral atoms. Our demonstration in a tweezer clock platform 12, 13, 14, 15, 9, 16 enables a circuit-based approach to quantum metrology with neutral atom optical clocks. To this end, we demonstrate two-qubit entangling gates with $99.62\%$ fidelity via Rydberg interactions 17, 18, 15 and dynamical connectivity 19 for optical clock qubits, which we combine with local addressing 16 to implement universally programmable quantum circuits. Using this approach, we generate a near-optimal entangled probe state 1, 4, a cascade of Greenberger-Horne-Zeilinger (GHZ) states of different sizes, and perform dual-quadrature 5 GHZ readout. We also show repeated fast phase detection with non-destructive conditional reset of clock qubits and minimal dead time between repetitions by implementing ancilla-based quantum logic spectroscopy 20 (QLS) for neutral atoms. Finally, we extend this to multi-qubit parity checks and measurement-based, heralded, Bell state preparation 21, 22, 23, 24. Our work lays the foundation for hybrid processor-clock devices with neutral atoms and more generally points to a future of practical applications for quantum processors linked with quantum sensors25.
@@
To demonstrate such complex entangled probe-state generation and repeated ancilla-based readout, we realize a scalable universal quantum processor with neutral atom qubits encoded in an ultranarrow optical transition (Fig. 1a). Importantly, the same experimental system can be utilized as a neutral atom optical clock 8, which currently forms the most stable frequency reference 35, 36. Our system specifically is a tweezer optical clock 12, 13, 14, 15, 9, 16, composed of neutral strontium-88 trapped in a tweezer array 42, 43 (Fig. 1b,c), which we now equip with universal quantum computing and repeated ancilla-based readout capabilities.
In particular, we demonstrate high-fidelity entangling gates 27, 17, 18 for optical clock qubits 15, achieving a state-of-the-art controlled-Z (CZ) fidelity for neutral atoms of $0.9962(4)$ (averaged over two-qubit symmetric stabilzier states), and generate an array of optical-clock-transition Bell pairs with a record fidelity. We further show that optical coherence is preserved during qubit transport in tweezers, previously realized with hyperfine qubits 19, 44, enabling dynamical circuit reconfiguration. Combined with single-site control 16 and mid-circuit readout 45, this allows us to take a circuit-based approach to realize strategies for entangled state preparation and readout designed to improve metrology (Fig. 1c).
High-fidelity entangling gates for optical qubits We begin by demonstrating high-fidelity entangling gates for the ultranarrow optical clock transition. The optical transition defines a qubit composed of the ground state of strontium-88 $\mathrm{{}^{1}S_{0}}$, denoted as $|0\rangle$, and the metastable state $\mathrm{{}^{3}P_{0}}$, denoted as $|1\rangle$ (Fig. 1b). To realize high-fidelity entangling operation, we apply a single phase-modulated Rydberg pulse for a CZ gate, which was shown to be time-optimal 46, 47 and recently implemented in hyperfine and nuclear atomic qubits 17, 18. To simultaneously realize high-fidelity single-qubit rotations, we operate at a magnetic field of 450 G which balances between sufficiently high Rabi frequency on the clock transition (scaling linearly with the magnetic field 48) and sufficiently strong interactions (as the Rydberg interaction strength is magnetic field dependent, see Methods, Ext. Data Fig. 1). We prepare atoms in the $|1\rangle$ state and perform erasure-cooling 45 which prepares atoms close to the motional ground state.
@@
We characterize the performance of such gates via a randomized circuit (Fig. 2a) designed to benchmark the CZ gate fidelity averaged over the symmetric stabilizer states, which are the set of states accessible with global single-qubit Clifford gates and two-qubit CZ gates (Methods, Ext. Data Fig. 3). Varying the number of CZ gates $\mathrm{N_{CZ}}$, we infer a CZ fidelity of $0.9962(4)$ by fitting the decay of the return probability as a function of $\mathrm{N_{CZ}}$ to an exponential trend (Fig. 2b). The quoted fidelity is corrected for false contribution from leakage out of the qubit manifold during the gate (Methods). We numerically find this gate benchmark to be within $5\%$ of the infidelity averaged over Haar random states for our gate parameters. Our result sets a new state-of-the-art fidelity in Rydberg-based two-qubit entangling gates for long-lived qubits 17, 49.
As an important application of the CZ gate, we generate an array of Bell pairs with optical clock qubits (Fig. 2c). After initializing atoms in superposition states and applying a CZ-gate, a $\pi/4$ pulse rotates the state into the even-parity, metrologically relevant, Bell state $|\Phi^{+}\rangle=(1/\sqrt{2})(|00\rangle+|11\rangle)$. For the state obtained in the experiment, we measure the population overlap with the ideal Bell state as well as drive parity oscillations (Fig. 2d) to infer the off-diagonal (coherence) terms 50. Overall, we obtain a Bell state generation fidelity of $0.976^{+4}_{-6}$ ($0.989^{+4}_{-6}$ SPAM corrected) which is limited, beyond the CZ fidelity, by the finite fidelity of clock $\pi/2$ pulses and decoherence during the circuit (Methods).
@@
Having demonstrated high-fidelity gates on optical clock qubits, we now turn to preparing larger entangled states and executing algorithms requiring more complex circuits. Specifically, we first demonstrate the preparation of a complex quantum register with metrological significance based on a cascade of entangled GHZ states. Such GHZ states enable, in principle, the best sensitivity for linear observables allowed by quantum theory - the Heisenberg limit - with precision scaling as $1/N$, where $N$ is the atom number. However, for a single $N$-qubit GHZ state, while the sensitivity to phase variations is increased by a factor of $N$, the dynamic range is likewise reduced by a factor of $N$ 3, 4, 37. A similar trade-off exists for spin-squeezed entangled states 38, 39.
Specifically, in current neutral atom optical atomic clocks, a smaller dynamic range also implies effectively reduced clock stability. Dynamic range here refers to the maximum interval over which phase excursions of a clock laser (the local oscillator) can be unambiguously measured and corrected for. Since phase excursions grow on average with time, a smaller dynamic range means a shorter allowable interrogation time. As the precision of phase estimation is reduced with shorter interrogation time, a smaller dynamic range results in a loss of precision. Therefore, in practice, the nominally improved sensitivity of GHZ states is counteracted by the loss of dynamic range, and no improvement of clock stability from fixed-$N$ GHZ states is expected over optimal interrogation with uncorrelated atoms 3, 1, 4, 37. For spin-squeezed states, similarly, only moderate improvements are predicted 38, 39, 7.
@@
A proposed scheme 1, 4 to circumvent this trade-off between sensitivity and dynamic range requires the simultaneous preparation of a cascade of $M$ different groups of GHZ-like states of the form $|\psi_{j}\rangle\propto(|0\rangle^{\otimes 2^{j}}+|1\rangle^{\otimes 2^{j}})$ with $j=0,1...M-1$. Such a probe state enables Heisenberg-limited sensitivity, up to logarithmic corrections, while preserving the dynamic range. It thus enables a viable path to quantum-enhanced metrology under realistic conditions where local oscillator noise is the main limitation, as in current neutral atom optical clocks 51, 37. We note that these proposals are quantum-enhanced extensions of the classical optimal interrogation schemes 5, 52 that were recently experimentally shown 16, 53.
@@
We consider different circuit realizations for generating the required simultaneous preparation of GHZ states with $j=0,1,2$, utilizing our universal quantum processor aided by dynamical array reconfiguration. While large GHZ states can be prepared with circuits applying extensive local control (Ext. Data Fig. 5), here we realize a circuit compiled with global control up to its final layer (Fig. 3a). This choice is guided by minimizing idle time added by local operations, during which laser-atom dephasing occurs. The local control for single-qubit rotations is achieved through local phase shifts imparted via sub-wavelength shifts in atomic position 16, and the connectivity for two-qubit gates is realized by mid-circuit array reconfiguration (Fig. 3b). We verify the simultaneous preparation of increasingly phase-sensitive states by varying the phase of a final $\pi/2$ pulse and measuring the parity signal $\mathcal{P}=\prod_{i}{Z_{i}}$ for each GHZ state. In Fig. 3c, we plot this signal against phase and find a 1:$1.96(6)$:$3.9(1)$ ratio in phase sensitivity, consistent with the 1:2:4 ratio expected.
@@
This demonstration measures only one quadrature of the optical phase signal. Importantly, the dynamic range can be doubled if both phase quadratures are probed simultaneously, so-called dual-quadrature readout 4, 5, 16, 53. We demonstrate such dual-quadrature readout for entangled states (Fig. 3d) by preparing two 4-atom GHZ states and then imparting a phase rotation to only one through local phase shifts. This enables readout in the ${\hat{Y}}$-basis in parallel to readout in the ${\hat{X}}$-basis. Here, a $\pi/8$ phase shift per atom ($\pi/2$ phase on the entangled 4-atom register) is realized by moving 16 all four atoms by $\lambda/16\approx 44$ nm. We note that such phase shifts could alternatively be realized by moving a single atom in the GHZ state by $\lambda/4$.
@@
An important question concerns the limitations on the maximum attainable GHZ state size in terms of achievable fidelity. We hypothesize that the phase noise of our clock laser is the dominant error source, which we test with experimental and numerical investigations. In the simultaneous GHZ state preparation scheme, we achieve a bare parity contrast of $0.52(3)$ for $N$=4 (Fig. 3c). To reduce the impact of clock laser phase noise, we prepare $N$=4 and $N$=2 GHZ states using a circuit with shorter idle time, i.e. wait time during which operations are done on other registers (by removing the single site operation $Z^{\frac{1}{4}}$ shown in Fig. 3a). With this sequence, we find an improved contrast of $0.68(3)$ and fidelity of $0.71(2)$ (Ext. Data Fig. 9). We note that the contrasts from both sequences are at or above the threshold of $0.51/\sqrt{N}=$ 0.5 required for a metrological gain 54, 2 at a fixed, short interrogation time. However, they do not exceed the higher threshold for the cascaded-GHZ scheme (Methods).
@@
Motivated by this, we study fidelity as a function of the circuit idle time. We measure a decay in fidelity and a modification of the bitstring distribution (Ext. Data Fig. 9) which follows the prediction of our error model, consistent with our hypothesis that clock laser phase noise is the dominant error source. We further run a model using a frequency noise spectrum of a 26-mHz clock laser system 55, for which we predict a $N$=4 GHZ state fidelity $\approx$ 0.97 (Methods, Ext. Data Fig. 6). Thus, we expect significantly improved results with an upgraded clock laser system, surpassing the metrological gain threshold for the cascaded-GHZ scheme, even for larger $N$ (Methods, Ext. Data Fig. 7).
@@
Figure 4: Repeated ancilla-based quantum logic spectroscopy (QLS). a, We consider three ensembles of atoms: reference atoms (R; red), clock atoms (C; purple), and ancillas (A; blue). An entangling gate maps the state of clock atoms to the state of ancilla atoms before mid-circuit readout (MCR) of the ancilla state. b, Logic spectroscopy via ancilla atoms. A Ramsey interrogation is performed in parallel on reference atoms (R) and clock atoms (C). A clock-ancilla CNOT gate maps the state onto ancilla atoms (A) for readout. c, Ancilla-based $\hat{X}$-measurement. A non-destructive measurement of the state $|\psi\rangle_{C}$ is realized by shelving clock qubits ($S$ operation) followed by ancilla MCR. We show that following unshelving ($S^{\dagger}$), the state of the clock qubit is projected to a well-defined state in the measurement basis, conditioned on the ancilla state. d, $\rm{N_{cyc}}$ repeated rounds of QLS with minimal (2.9 ms) dead time, enabled by ancilla replacement via array reconfiguration. Each round $j$ consists of a fixed evolution time $T_{j}$ followed by ancilla-based $\hat{X}$ measurement before a final round where the clock qubit is directly measured in a Ramsey sequence with variable time t. e, We find improved contrast in the final measurement round when conditioned on ancilla results (see text) in all previous rounds (purple markers) as compared to reference clock atoms which are protected during MCR but do not interact with ancillas (red markers).
@@
To efficiently realize ancilla-based readout in atomic systems, information-carrying data qubits must first be protected from global imaging light and their coherence must be preserved during mid-circuit readout (MCR) of ancilla qubits. Such protection can be realized by using dual-species arrays 56, cavity-assisted readout 57, a separate readout zone 24, or shelving to states dark to imaging light 58, 59, 60. In our system, we take the last approach and have recently realized MCR 45 by shelving optical qubits into superpositions of motional states which are long-lived and dark to imaging light.
@@
To realize a non-destructive ancilla-based $\hat{X}$-measurement, we then shelve clock qubits and read out the ancilla state. To this end, we ramp up the depth of ancilla traps by a factor of 5 (leaving traps with clock atoms unchanged) such that they are non-resonant with the global motional shelving drive. Following shelving of clock qubits, we perform ancilla MCR of the signal field $\Delta$. Conditioned on the ancilla measurement result, the clock atom is projected to a known state in the measurement basis, which enables repeated interrogation. We verify this by measuring the clock qubit state in the $\hat{X}$-basis via a $\pi/2$ pulse with a variable phase $\phi$ before detection (Fig. 4c). We find the clock qubit is projected onto the |+ket|$\rangle$| + ⟩ (|ket|$\rangle$| - ⟩) state conditioned on the ancilla measurement result $|0\rangle$ ($|1\rangle$) in the $\hat{Z}$-basis, with a contrast of $0.84(4)$ ($0.73(4)$). These results are conditioned on successful one-way shelving 45, 18 (Methods).
@@
Ideally, non-destructive readout enables repetitive measurements of the same qubit or the same register. This can be achieved by resetting the ancilla qubit and repeating the readout cycle. In our case, we utilize an 18 $\mu$s fast imaging method 61 which is much shorter than the qubit coherence time but results in a high probability of losing the ancilla atom during imaging. We thus realize repeated rounds of measurement and reset by replacing the ancilla atom with a reserve atom, which was shelved during ancilla MCR, using dynamical reconfiguration. This yields a dead time of only 2.9 ms dominated by shelving and unshelving operations. Alternatively, we demonstrate reusing the same ancilla atom by applying a 10-ms low-loss imaging followed by re-cooling (Methods, Ext. Data Fig. 10).
@@
We now turn to demonstrate another relevant application of the conditional parity projection in a weight-2 ancilla MCR: long-range entanglement induced by measurement 21, 22, 23, 24. We first prepare a 3-qubit cluster state and then perform MCR of the central qubit in the $\hat{X}$-basis (Fig. 5c). As $\hat{Z}\hat{X}\hat{Z}$ is a stabilizer of the cluster state, this measurement generates a Bell pair with a well-defined parity for the exterior atoms. Conditioning on the ancilla result $|0\rangle$, we find a Bell state with population $0.71(2)$ in the even parity subspace $\{|00\rangle,|11\rangle\}$ and a parity oscillation contrast of $0.52(3)$, yielding an overall raw fidelity of $0.61(2)$ (Fig. 5d). If instead we condition on the ancilla result $|1\rangle$, we find the state is projected to the odd parity Bell state, which is not useful in a metrological context, with $0.64(2)$ in the odd parity subspace $\{|01\rangle,|10\rangle\}$. The scheme demonstrated here can be considered a heralded preparation, which could be made deterministic by a feedback operation 21, 22, 23, 24. We note that with improvements in fidelity, and with sufficiently fast feedback, this could then be readily extended to perform measurement-based preparation of large-scale GHZ states with only a constant depth circuit 22, 21, 23. This is in contrast to linear or log-depth standard unitary circuits 31, which could be instrumental for generating large-scale optical clock GHZ states in future devices.
@@
  Our results demonstrate how universal quantum processing and ancilla-based readout can be integrated with a neutral atom optical clock. Extending such schemes to larger two-dimensional tweezer clock systems 14 would enable the creation of multiple copies of GHZ states at a given size. We note in particular that the schemes demonstrated here could be easily generalized to two dimensions by performing all operations in parallel for a set of rows to generate multiple copies of GHZ states. We see this as a viable avenue towards quantum-enhanced operation of neutral atom clocks using cascades of GHZ states 1, 4 combined with classically optimal interrogation schemes and quadrature readout 5, 52, 16, 53.
@@
We have demonstrated a CZ fidelity of $0.9962(4)$ , averaged over the two-qubit symmetric stabilizer states, setting a new state-of-the-art for entangling gate fidelities in long-lived states of neutral atoms, as well as the full toolbox needed to execute relevant algorithms for metrology and computation with optical clock qubits. The main limitation to current circuit fidelities is the clock laser phase noise resulting in dephasing during circuit idle times. The ancilla-based MCR is further limited by the fidelity of shelving qubits in motional states. Upgrading the clock laser system used in this work to match the current state-of-the-art 35, 14 would enable generating significantly larger GHZ states than demonstrated here (Ext. Data Fig. 6, Methods). We then expect the maximum GHZ state size will eventually be limited by a combination of clock laser phase noise and CZ gate fidelity. In this context, we note that the CZ gate fidelity is currently dominated by phase noise of the Rydberg laser (Ext. Data Fig. 4, Methods), which could easily be improved by implementing cavity filtering techniques 62. Further, utilizing native, multi-qubit gates 63, 17 or effective all-to-all connectivity 19 could reduce the required circuit depth. Ancilla-based MCR would also improve with a higher stability clock laser and better control and homogenization of trap waists 45. Finally, GHZ states can be prepared with constant depth circuits via measurement-based schemes 21, 22, for which we showed first steps with optical clock qubits in Fig. 5c,d.

### v3

@@
Enhancing the precision of measurements by harnessing entanglement is a long-sought goal in the field of quantum metrology 1, 2. Yet attaining the best sensitivity allowed by quantum theory in the presence of noise is an outstanding challenge, requiring optimal probe-state generation and readout strategies 3, 4, 5, 6, 7. Neutral atom optical clocks 8, leading systems for measuring time, have shown recent progress in terms of entanglement generation 9, 10, 11, but currently lack the control capabilities to realize such schemes. Here we show universal quantum operations and ancilla-based readout for ultranarrow optical transitions of neutral atoms. Our demonstration in a tweezer clock platform 12, 13, 14, 15, 9, 16 enables a circuit-based approach to quantum metrology with neutral atom optical clocks. To this end, we demonstrate two-qubit entangling gates with $99.62(3)\%$ fidelity – averaged over symmetric input states – via Rydberg interactions 17, 18, 15 and dynamical connectivity 19 for optical clock qubits, which we combine with local addressing 16 to implement universally programmable quantum circuits. Using this approach, we generate a near-optimal entangled probe state 1, 4, a cascade of Greenberger-Horne-Zeilinger (GHZ) states of different sizes, and perform dual-quadrature 5 GHZ readout. We also show repeated fast phase detection with non-destructive conditional reset of clock qubits and minimal dead time between repetitions by implementing ancilla-based quantum logic spectroscopy 20 (QLS) for neutral atoms. Finally, we extend this to multi-qubit parity checks and measurement-based, heralded, Bell state preparation 21, 22, 23, 24. Our work lays the foundation for hybrid processor-clock devices with neutral atoms and more generally points to a future of practical applications for quantum processors linked with quantum sensors25.
@@
To demonstrate such complex entangled probe-state generation and repeated ancilla-based readout, we realize a scalable universal quantum processor with neutral atom qubits encoded in an ultranarrow optical transition (Fig. 1a). Importantly, the same experimental system can be utilized as a neutral atom optical clock 8, which currently forms the most stable frequency reference 35, 36. Our system specifically is a tweezer optical clock 12, 13, 14, 15, 9, 16, composed of neutral strontium-88 trapped in a tweezer array 42 (Fig. 1b,c), which we now equip with universal quantum computing and repeated ancilla-based readout capabilities.
In particular, we demonstrate high-fidelity entangling gates 27, 17, 18 for optical clock qubits 15, achieving a state-of-the-art controlled-Z (CZ) fidelity for neutral atoms of $0.9962(3)$ (averaged over two-qubit symmetric states), and generate an array of optical-clock-transition Bell pairs with a record fidelity. We further show that optical coherence is preserved during qubit transport in tweezers, previously realized with hyperfine qubits 19, 43, enabling dynamical circuit reconfiguration. Together with high-fidelity global single-qubit rotations, single-site control 16 and mid-circuit readout 44, this allows us to take a circuit-based approach to realize strategies for entangled state preparation and readout designed to improve metrology (Fig. 1c).
High-fidelity entangling gates for optical qubits We begin by demonstrating high-fidelity entangling gates for the ultranarrow optical clock transition. The optical transition defines a qubit composed of the ground state of strontium-88 $\mathrm{{}^{1}S_{0}}$, denoted as $|0\rangle$, and the metastable state $\mathrm{{}^{3}P_{0}}$, denoted as $|1\rangle$ (Fig. 1b). To realize high-fidelity entangling operation, we apply a single phase-modulated Rydberg pulse for a CZ gate, which was shown to be time-optimal 45, 46 and recently implemented in hyperfine and nuclear atomic qubits 17, 18. To simultaneously realize high-fidelity single-qubit rotations, we operate at a magnetic field of 450 G which balances between sufficiently high Rabi frequency on the clock transition (scaling linearly with the magnetic field) and sufficiently strong interactions (as the Rydberg interaction strength is magnetic field dependent, see Methods, Ext. Data Fig. 1). We prepare atoms in the $|1\rangle$ state and perform erasure-cooling 44 which prepares atoms close to the motional ground state.
@@
We characterize the performance of such gates via a randomized circuit (Fig. 2a) designed to benchmark the CZ gate fidelity averaged over the two-qubit symmetric states (Methods, Ext. Data Fig. 3). Varying the number of CZ gates $\mathrm{N_{CZ}}$, we infer a CZ fidelity of $0.9962(3)$ by fitting the decay of the return probability as a function of $\mathrm{N_{CZ}}$ to an exponential trend (Fig. 2b). The quoted fidelity is corrected for false contribution from leakage out of the qubit manifold during the gate (Methods). We numerically find this gate benchmark to be within $5\%$ of the infidelity averaged over Haar random states for our gate parameters 47. Our result sets a new state-of-the-art fidelity in Rydberg-based two-qubit entangling gates for long-lived qubits 17, 48.
As an important application of the CZ gate, we generate an array of Bell pairs with optical clock qubits (Fig. 2c). After initializing atoms in superposition states and applying a CZ-gate, a $\pi/4$ pulse rotates the state into the even-parity, metrologically relevant, Bell state $|\Phi^{+}\rangle=(1/\sqrt{2})(|00\rangle+|11\rangle)$. For the state obtained in the experiment, we measure the population overlap with the ideal Bell state as well as drive parity oscillations (Fig. 2d) to infer the off-diagonal (coherence) terms 49. Overall, we obtain a Bell state generation fidelity of $0.976^{+4}_{-6}$ ($0.989^{+4}_{-6}$ SPAM corrected) which is limited, beyond the CZ fidelity, by the finite fidelity of clock $\pi/2$ pulses and decoherence during the circuit (Methods).
@@
Having demonstrated high-fidelity gates on optical clock qubits, we now turn to preparing larger entangled states and executing algorithms requiring more complex circuits. Specifically, we first demonstrate the preparation of a complex quantum register with metrological significance based on a cascade of entangled GHZ states. Such GHZ states enable, in principle, the best sensitivity for linear observables allowed by quantum theory - the Heisenberg limit - with precision scaling as $1/N$, where $N$ is the atom number. However, for a single $N$-qubit GHZ state, while the sensitivity to phase variations is increased by a factor of $N$, the dynamic range of phases that can be unambiguously measured is likewise reduced by a factor of $N$ 3, 4, 37. A similar trade-off exists for spin-squeezed entangled states 38, 39. Importantly, a smaller dynamic range means a shorter allowable interrogation time, resulting in precision loss.
@@
A proposed scheme 1, 4 to circumvent this trade-off between sensitivity and dynamic range requires the simultaneous preparation of a cascade of $M$ different groups of GHZ-like states of the form $|\psi_{j}\rangle\propto(|0\rangle^{\otimes 2^{j}}+|1\rangle^{\otimes 2^{j}})$ with $j=0,1...M-1$. Such a probe state enables Heisenberg-limited sensitivity, up to logarithmic corrections, while preserving the dynamic range. It thus enables a viable path to quantum-enhanced metrology under realistic conditions where local oscillator noise is the main limitation, as in current neutral atom optical clocks 37. We note that these proposals are quantum-enhanced extensions of the classical optimal interrogation schemes 5, 50 that were recently experimentally shown 16, 51.
@@
We consider different circuit realizations for generating the required simultaneous preparation of GHZ states with $j=0,1,2$, utilizing our universal quantum processor aided by dynamical array reconfiguration. While large GHZ states can be prepared with circuits applying extensive local control (Ext. Data Fig. 5), here we realize a circuit compiled with global control up to its final layer (Fig. 3a). This choice is guided by minimizing idle time added by local operations, during which laser-atom dephasing occurs. The local control for single-qubit rotations is achieved through local phase shifts imparted via sub-wavelength shifts in atomic position 16, and the connectivity for two-qubit gates is realized by mid-circuit array reconfiguration (Fig. 3b). We verify the simultaneous preparation of increasingly phase-sensitive states by varying the phase of a final $\pi/2$ pulse and measuring the parity signal $\mathcal{\hat{P}}=\prod_{i}{\hat{Z}_{i}}$ for each GHZ state. In Fig. 3c, we plot this signal against phase and find a 1:$1.96(6)$:$3.9(1)$ ratio in phase sensitivity, consistent with the 1:2:4 ratio expected.
@@
This demonstration measures only one quadrature of the optical phase signal. Importantly, the dynamic range can be doubled if both phase quadratures are probed simultaneously, so-called dual-quadrature readout 4, 5, 16, 51. We demonstrate such dual-quadrature readout for entangled states (Fig. 3d) by preparing two 4-atom GHZ states and then imparting a phase rotation to only one through local phase shifts. This enables readout in the ${\hat{Y}}$-basis in parallel to readout in the ${\hat{X}}$-basis. Here, a $\pi/8$ phase shift per atom ($\pi/2$ phase on the entangled 4-atom register) is realized by moving 16 all four atoms by $\lambda/16\approx 44$ nm. We note that such phase shifts could alternatively be realized by moving a single atom in the GHZ state by $\lambda/4$.
@@
An important question concerns the limitations on the maximum attainable GHZ state size in terms of achievable fidelity. We hypothesize that the phase noise of our clock laser is the dominant error source, which we test with experimental and numerical investigations. In the simultaneous GHZ state preparation scheme, we achieve a bare parity contrast of $0.52(3)$ for $N$=4 (Fig. 3c). To reduce the impact of clock laser phase noise, we prepare $N$=4 and $N$=2 GHZ states using a circuit with shorter idle time, i.e. wait time during which operations are done on other registers (by removing the single site operation $\hat{Z}^{\frac{1}{4}}$ shown in Fig. 3a). With this sequence, we find an improved contrast of $0.68(3)$ and fidelity of $0.71(2)$ (Ext. Data Fig. 9). We note that the contrasts from both sequences are at or above the threshold of $0.51/\sqrt{N}=$ 0.5 required for a metrological gain 52, 2 at a fixed, short interrogation time. However, they do not exceed the higher threshold for the cascaded-GHZ scheme (Methods).
@@
Motivated by this, we study fidelity as a function of the circuit idle time. We measure a decay in fidelity and a modification of the bitstring distribution (Ext. Data Fig. 9) which follows the prediction of our error model, consistent with our hypothesis that clock laser phase noise is the dominant error source. We further run a model using a frequency noise spectrum of a 26-mHz clock laser system 53, for which we predict a $N$=4 GHZ state fidelity $\approx$ 0.97 (Methods, Ext. Data Fig. 6). Thus, we expect significantly improved results with an upgraded clock laser system, surpassing the metrological gain threshold for the cascaded-GHZ scheme, even for larger $N$ (Methods, Ext. Data Fig. 7).
@@
Figure 4: Repeated ancilla-based quantum logic spectroscopy (QLS). a, We consider three ensembles of atoms: reference atoms (R; red), clock atoms (C; purple), and ancillas (A; blue). An entangling gate maps the state of clock atoms to the state of ancilla atoms before mid-circuit readout (MCR) of the ancilla state. b, Logic spectroscopy via ancilla atoms. A Ramsey interrogation is performed in parallel on reference atoms (R) and clock atoms (C). A clock-ancilla CNOT gate (controlled on qubit state $|0\rangle$) maps the state onto ancilla atoms (A) for readout. c, Ancilla-based $\hat{X}$-measurement. A non-destructive measurement of the state $|\psi\rangle_{C}$ is realized by shelving clock qubits ($\hat{S}$ operation) followed by ancilla MCR. We show that following unshelving ($\hat{S}^{\dagger}$), the state of the clock qubit is projected to a well-defined state in the measurement basis, conditioned on the ancilla state. d, $\rm{N_{cyc}}$ repeated rounds of QLS with minimal (2.9 ms) dead time, enabled by ancilla replacement via array reconfiguration. Each round $j$ consists of a fixed evolution time $T_{j}$ followed by ancilla-based $\hat{X}$ measurement before a final round where the clock qubit is directly measured in a Ramsey sequence with variable time t. e, We find improved contrast in the final measurement round when conditioned on ancilla results (see text) in all previous rounds (purple markers) as compared to reference clock atoms which are protected during MCR but do not interact with ancillas (red markers).
@@
To efficiently realize ancilla-based readout in atomic systems, information-carrying data qubits must first be protected from global imaging light and their coherence must be preserved during mid-circuit readout (MCR) of ancilla qubits. Such protection can be realized by using dual-species arrays 40, cavity-assisted readout 54, a separate readout zone 24, or shelving to states dark to imaging light 55, 56, 57. In our system, we take the last approach and have recently realized MCR 44 by shelving optical qubits into superpositions of motional states which are long-lived and dark to imaging light.
@@
To realize a non-destructive ancilla-based $\hat{X}$-measurement, we then shelve clock qubits and read out the ancilla state. To this end, we ramp up the depth of ancilla traps by a factor of 5 (leaving traps with clock atoms unchanged) such that they are non-resonant with the global motional shelving drive. Following shelving of clock qubits, we perform ancilla MCR of the signal field $\Delta$. Conditioned on the ancilla measurement result, the clock atom is projected to a known state in the measurement basis, which enables repeated interrogation. We verify this by measuring the clock qubit state in the $\hat{X}$-basis via a $\pi/2$ pulse with a variable phase $\phi$ before detection (Fig. 4c). We find the clock qubit is projected onto the |+ket|$\rangle$| + ⟩ (|ket|$\rangle$| - ⟩) state conditioned on the ancilla measurement result $|0\rangle$ ($|1\rangle$) in the $\hat{Z}$-basis, with a contrast of $0.84(4)$ ($0.73(4)$). These results are conditioned on successful one-way shelving 44, 18 (Methods).
@@
Ideally, non-destructive readout enables repetitive measurements of the same qubit or the same register. This can be achieved by resetting the ancilla qubit and repeating the readout cycle. In our case, we utilize an 18 $\mu$s fast imaging method 58 which is much shorter than the qubit coherence time but results in a high probability of losing the ancilla atom during imaging. We thus realize repeated rounds of measurement and reset by replacing the ancilla atom with a reserve atom, which was shelved during ancilla MCR, using dynamical reconfiguration. This yields a dead time of only 2.9 ms dominated by shelving and unshelving operations. Alternatively, we demonstrate reusing the same ancilla atom by applying a 10-ms low-loss imaging followed by re-cooling (Methods, Ext. Data Fig. 10).
@@
We now turn to demonstrate another relevant application of the conditional parity projection in a weight-2 ancilla MCR: long-range entanglement induced by measurement 21, 22, 23, 24. We first prepare a 3-qubit cluster state and then perform MCR of the central qubit in the $\hat{X}$-basis (Fig. 5c). As $\hat{Z}\hat{X}\hat{Z}$ is a stabilizer of the cluster state, this measurement generates a Bell pair with a well-defined parity for the exterior atoms. Conditioning on the ancilla result $|0\rangle$, we find the Bell state $|\Phi^{+}\rangle$ with population $0.71(2)$ in the even parity subspace $\{|00\rangle,|11\rangle\}$ and a parity oscillation contrast of $0.52(3)$, yielding an overall raw fidelity of $0.61(2)$ (Fig. 5d). If instead we condition on the ancilla result $|1\rangle$, we find the state is projected to the odd parity Bell state $|\Psi^{-}\rangle=(1/\sqrt{2})(|01\rangle-|10\rangle)$, which is not metrologically useful in our context, with $0.64(2)$ in the odd parity subspace $\{|01\rangle,|10\rangle\}$. Alternatively, by performing an ancilla-based $\hat{X}\hat{X}$ measurement, one can prepare both metrologically relevant Bell states $|\Phi^{+}\rangle$,$|\Phi^{-}\rangle=(1/\sqrt{2})(|00\rangle-|11\rangle)$. Such schemes can be considered a heralded preparation, which could be made deterministic by a feedback operation 21, 22, 23, 24. We note that with improvements in fidelity, and with sufficiently fast feedback, this could then be readily extended to perform measurement-based preparation of large-scale GHZ states with only a constant depth circuit 22, 21, 23. This is in contrast to linear or log-depth standard unitary circuits 31, which could be instrumental for generating large-scale optical clock GHZ states in future devices.
@@
  Our results demonstrate how universal quantum processing and ancilla-based readout can be integrated with a neutral atom optical clock. Extending such schemes to larger two-dimensional tweezer clock systems 14 would enable the creation of multiple copies of GHZ states at a given size. We note in particular that the schemes demonstrated here could be easily generalized to two dimensions by performing all operations in parallel for a set of rows to generate multiple copies of GHZ states. We see this as a viable avenue towards quantum-enhanced operation of neutral atom clocks using cascades of GHZ states 1, 4 combined with classically optimal interrogation schemes and quadrature readout 5, 50, 16, 51.
@@
We have demonstrated a CZ fidelity of $0.9962(3)$ , averaged over the two-qubit symmetric states, setting a new state-of-the-art for entangling gate fidelity, as well as the full toolbox needed to execute relevant algorithms for metrology and computation with optical clock qubits. The main limitation to current circuit fidelities is the clock laser phase noise resulting in dephasing during circuit idle times. The ancilla-based MCR is further limited by the fidelity of shelving qubits in motional states. Upgrading the clock laser system used in this work to match the current state-of-the-art 35, 14 would enable generating significantly larger GHZ states than demonstrated here (Ext. Data Fig. 6, Methods). We then expect the maximum GHZ state size will eventually be limited by a combination of clock laser phase noise and CZ gate fidelity. In this context, we note that the CZ gate fidelity is currently dominated by phase noise of the Rydberg laser (Ext. Data Fig. 4, Methods), which could easily be improved by implementing cavity filtering techniques. Further, utilizing native, multi-qubit gates 59, 17 or effective all-to-all connectivity 19 could reduce the required circuit depth. Ancilla-based MCR would also improve with a higher stability clock laser and better control and homogenization of trap waists 44. Finally, GHZ states can be prepared with constant depth circuits via measurement-based schemes 21, 22, for which we showed first steps with optical clock qubits in Fig. 5c,d.